# Enabling Precise Topic Alignment in Large Language Models via Sparse Autoencoders

### v1

@@ 4.2 Layer-Level Output
Figure 5: (a) shows that Clamp is not sensitive to incoming token context. While we can observe in (b) that Swap varies the number of neurons changed, which reflects that it accounts for token context and feasible alignment potential.
@@
(a) Metrics for Clamp Approach.
(b) Metrics for Swap Approach.
Figure 6: Generally, the difference in reconstruction error for the modifications is less for aligned tokens than unaligned tokens. In Swap (b), we see lower contamination as a metric of uncertainty than with Clamp (a), which means that the neurons firing have higher alignment scores with Swap than Clamp. Special tokens not visualized.
In Fig. 6, the unaligned text has a higher difference in reconstruction error because the the more generic SAE trained to reduce reconstruction error should better represent the unaligned text while the SAE modifications should better represent the aligned text. We have excluded special tokens from our analysis, which had very high reconstruction error with our modifications. We hypothesize this is because we removed these special tokens during our scoring process, but they were included while training the underlying SAEs. This could be why the tokens with the lowest/highest differences in reconstruction error are the same between both mechanisms. Still, in the Swap approach, we tend to notice lower contamination (x-axis in Fig. 6(b)), which is in line with our expectations that Swap is better able to align layer-level output based on our alignment scores.
@@ 4.3 Generated Output
Example Text: My heart is broken, so I should laryngosc- -opy to visualize my heart. Used cricoid balloon tubes to visualize the chest. I also used cricoid balloon tubes to visualize my stomach. so thanking my friends. My heart breaks so my heart breaks 0 [Good] just tell my husband and wife to get out. [..] My kids have a daughter and she is a big sister and I am so jealous I will take a divorce. be n,.) n ns- is, = ] is. n and] is] on is ] =] n is ] is a) n] to get [ so. so [ is.] n].] and was a lot] on the] in the in ons and]] ns’] should go to sleep. It’s not liek my brain is doing any work t oget me back up and working again. My head has stopped
6 [Poor] I should be, I was-c-c ””R and then he was the C-I was-r, I was R R (I was) n I-P R (The ””R R R-T) The first is-C) ?I I-3-in–, it? I… ””L?, ””L, ””……,, ””I?,, ””A I I I?…, I-A?, I I?, ””A?I-A……,I-I,,,? not the ””c and I have it-f, the ””x is a very-al ”” ””- and it-r-R R and a very) R-. (I was not the two)
@@
Table 2: Standard metrics in SAE layers with different representativeness (0 is better than 6 see Fig. 3). Our results show that the standard perplexity metric is flawed (high values for FT in the medical domain), the MI baseline consistently returns outputs with low linguistic meaning, the swap approach is competitive on distance to alignment text (see Appendix for more details), and the contamination score determined by the topic alignment process aligns with the output quality on inspection.
We compare our approach to an out-of-the-box fine-tuning method, adapted from HuggingFace (a), and generated 64 tokens per prompt with a top-k value of 5. To demonstrate alignment power, our input prompts were common medical metaphors (see Appendix for full prompts), as we expect that the general-purpose LLM would generate text related to the metaphor, while the alignment approaches should generate medical text. Our results in Table 4.3 show that approach performance depends on the underlying SAE’s representational power:
@@ 5 Discussion and Conclusion
Further, research into improving the representational power of the underlying SAEs, testing on rare alignment topics, and studying patterns of coactivation can further the potential of using SAEs for topic alignment as an alternative to fine tuning. Finally, it is worth noting that this approach can also address other limitations with fine-tuning (e.g., where privacy is essential, as ${\mathbb{P}}_{align}$ tokens are never directly used for scoring, or when there are too few examples for successful fine tuning) and warrants exploration on these dimensions as well. .
@@ A.2 Design Choices for Scoring Configurations
Figure 10: Scores across different listed configurations
Choice of $dist$: For four topics from the P3 set relatively individual tasks (Amazon, Yelp, IMDB, and Wikipedia), we sampled 20 prompts and calculated inter and intra-Euclidian distance metrics as shown in Table 5.
Task Distances
Intra-topic Inter-topic
Yelp 1.357 $\pm$ 0.005 1.128$\pm$ 0.04
Wiki 1.412 $\pm$ 0.004 1.144 $\pm$ 0.041
Amazon 1.34 $\pm$ 0.006 0.95 $\pm$ 0.04
IMDB 1.34 $\pm$ 0.005 1.00 0.04
Table 5: Distance Metric Evaluation
Task Text Excerpt Compare Inter-Topic Excerpt
Amazon ” I purchased these deck of cards[…]layout or spreads are too big, and it comes with a lil booketlet not really a booklet […] do a reading for a man and women […] Very disappointed” Yelp ”I used to go here for tech books a lot. I went in and the section for tech books is half the size it used to be, completely disheveled, and contains no organization beyond the dummies books”
Amazon ” I guess there are times when the majority of the reviews do not live up to the movie. This movie was horrible to say the least. Horrible plot, acting, and props. Don’t waste your time on this on, not even if you watch it for free.” IMDB ”[..]This is one of the most pointless films ever made. […] Its a wonder that it was ever put on video.[..] Surely this film is a waste of the money used to create it, and a waste of anyone’s time watching it.”
Amazon ”” I guess there are times when the majority of the reviews do not live up to the movie. This movie was horrible to say the least. Horrible plot, acting, and props. Don’t waste your time on this on, not even if you watch it for free.” IMDB ”Can only be described as awful. […] impossible to believe that it can get worse - but fear not because it does.[…] script is worse than the acting or whether the directing is worse than both. […] Lucky it only cost me \$1 to hire.
Amazon ”This is a good read. A little slow at times. I recommend the entire Tucker Mills series. Yelp ”ALWAYS very consistent. Food is ALWAYS good […] So yummy” ”
Table 6: Distance Examples
@@ A.3 Scoring Output
We found the distance metric reasonable because the intra-topic distance was less than the inter-topic distance per topic. Still, some examples of prompts from different topics had a distance less than the maximum distance between intra-topic prompts. Upon inspection, these distances still made sense because even though they were from different tasks, they discussed similar topics, as shown in Table 6.
@@ A.5 Generated Output
We provide additional correctness metrics for aligned prompts (from Gamino ) and unaligned prompts (from HuggingFace (b)) in Tables 7 and 8.
@@
Table 8: Results for different metrics across SAE layers
Finally, we show results for a LORA + Prompt tuning approach to supplement the standard fine-tuning approach (FT) in Tables 2-4:

### v2

@@ 4.2 Layer-Level Output
Figure 5: Clamp is not sensitive to incoming token context. We observe that Swap varies the number of neurons changed, which reflects that it accounts for token context and feasible alignment potential.
@@
Figure 6: Generally, the difference in reconstruction error for the modifications is less for aligned tokens than unaligned tokens. In Swap, we see lower contamination as a metric of uncertainty than with Clamp, which means that the neurons firing have higher alignment scores with Swap than Clamp. Special tokens not visualized.
In Fig. 6, the unaligned text has a higher difference in reconstruction error because the the more generic SAE trained to reduce reconstruction error should better represent the unaligned text while the SAE modifications should better represent the aligned text. We have excluded special tokens from our analysis, which had very high reconstruction error with our modifications. We hypothesize this is because we removed these special tokens during our scoring process, but they were included while training the underlying SAEs. This could be why the tokens with the lowest/highest differences in reconstruction error are the same between both mechanisms. Still, in the Swap approach, we tend to notice lower contamination (x-axis in Fig. LABEL:fig:subfig3s), which is in line with our expectations that Swap is better able to align layer-level output based on our alignment scores.
@@ 4.3 Generated Output
Example Text: My heart is broken, so I should laryngosc- -opy to visualize my heart. Used cricoid balloon tubes to visualize the chest. I also used cricoid balloon tubes to visualize my stomach. so thanking my friends. My heart breaks so my heart breaks 0 just tell my husband and wife to get out. [..] My kids have a daughter and she is a big sister and I am so jealous I will take a divorce. be n,.) n ns- is, = ] is. n and] is] on is ] =] n is ] is a) n] to get [ so. so [ is.] n].] and was a lot] on the] in the in ons and]] ns’] should go to sleep. It’s not liek my brain is doing any work t oget me back up and working again. My head has stopped
6 I should be, I was-c-c ””R and then he was the C-I was-r, I was R R (I was) n I-P R (The ””R R R-T) The first is-C) ?I I-3-in–, it? I… ””L?, ””L, ””……,, ””I?,, ””A I I I?…, I-A?, I I?, ””A?I-A……,I-I,,,? not the ””c and I have it-f, the ””x is a very-al ”” ””- and it-r-R R and a very) R-. (I was not the two)
@@
Table 2: Standard metrics in SAE layers with different representativeness (0 is more representative than 6 see Fig. 3). Our results show that the standard perplexity metric is flawed (high values for FT in the medical domain), the MI baseline consistently returns outputs with low linguistic meaning, the swap approach is competitive on distance to alignment text (see Appendix Tables 5 6 for more details), and the contamination score determined by the topic alignment process aligns with the output quality on inspection.
[…]
@@ 5 Discussion and Conclusion
Further, research into improving the representational power of the underlying SAEs, testing on rare alignment topics, and studying patterns of coactivation can further the potential of using SAEs for topic alignment as an alternative to fine tuning. Finally, it is worth noting that this approach can also address other limitations with fine-tuning (e.g., where privacy is essential, as ${\mathbb{P}}_{align}$ tokens are never directly used for scoring, or when there are too few examples for successful fine tuning) and warrants exploration on these dimensions as well.
@@ A.2 Design Choices for Scoring Configurations
Figure 10: Scores across different listed configurations.
@@ A.3 Scoring Output
We found the distance metric reasonable because the intra-topic distance was less than the inter-topic distance per topic. Still, some examples of prompts from different topics had a distance less than the maximum distance between intra-topic prompts. Upon inspection, these distances still made sense because even though they were from different tasks, they discussed similar topics, as shown in Table LABEL:distex.
@@ A.5 Generated Output
We provide additional correctness metrics for aligned prompts (from Gamino ) and unaligned prompts (from HuggingFace (b)) in Tables 5 and 6.
@@
Table 6: Results for different metrics across SAE layers in unaligned text show that Swap outperforms Clamp on Layer 0 (the more representative layer).
Finally, we show results for a LORA + Prompt tuning approach to supplement the standard fine-tuning approach (FT) in Tables 7: